OpenStudy (anonymous):

The radius of a clindrical gift box is (2x +3) inches. The height of the gift box is twice the radius. What is the surface area of the cylinder? Answer as polynomial in standard form.

OpenStudy (cggurumanjunath):

surface area of cylinder = ? @yourmomneedsanswers !

OpenStudy (cggurumanjunath):

https://www.google.co.in/#q=surface%20area%20of%20cylinder

OpenStudy (anonymous):

2πrh+2πr2, right?

OpenStudy (cggurumanjunath):

ya !

OpenStudy (anonymous):

so what do i do with that

OpenStudy (cggurumanjunath):

r= (2x+3) h=2(2x+3)

OpenStudy (cggurumanjunath):

h=2r ; given in the question !

OpenStudy (cggurumanjunath):

2pir(r+h)=2*pi*(2x+3)(2x+3+2(2x+3))

OpenStudy (cggurumanjunath):

sa = 2*pi*(2x+3)(2x+3+4x+6)=2*pi*(2x+3)(6x+9)

OpenStudy (cggurumanjunath):

do FOIL method and simplify !

OpenStudy (cggurumanjunath):

do u know FOIL method ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i mess up on it a lot

OpenStudy (cggurumanjunath):

ok.

OpenStudy (cggurumanjunath):

try it .

OpenStudy (cggurumanjunath):

i'll help u !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is so confusing

OpenStudy (cggurumanjunath):

try !

OpenStudy (cggurumanjunath):

(2x+3)(6x+9) = 2x(6x+9)+3(6x+9)

OpenStudy (cggurumanjunath):

=12x^2+18x+18x+27

OpenStudy (cggurumanjunath):

final answer =2*pi*(12x^2+36x+27)

OpenStudy (cggurumanjunath):

is that clear for you ? @yourmomneedsanswers

OpenStudy (anonymous):

that doesnt sound like the surface area, or maybe im not understanding?

OpenStudy (anonymous):

it says it isn't right

OpenStudy (cggurumanjunath):

24*pi*x^2+72*pi*x+54pi

OpenStudy (cggurumanjunath):

6*pi(4x^2+12x+9)

OpenStudy (anonymous):

that wasnt right

OpenStudy (cggurumanjunath):

then do u have the answer ?

OpenStudy (anonymous):

no.